OpenStudy (anonymous):

please help me find the integral of xlnx dx. i really would appreciate steps not just what rules to use. i know i am using integration by parts. i just get caught in a loop for some reason.

OpenStudy (jamesj):

So if you integrate once with u' = x and v = ln x, what do you get?

OpenStudy (anonymous):

i did u = lnx

OpenStudy (jamesj):

Use my suggestion

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so = x and dv = lnx is that how u say to start

OpenStudy (anonymous):

u = x*

OpenStudy (jamesj):

I'm saying u' = x and v = ln x

OpenStudy (anonymous):

if you are originally given xlnx, shouldnt derivative of u be 1

OpenStudy (jamesj):

The rule of integration by parts is \[ \int u' \ v \ dx \ = \ u \ v \ - \ \int u \ v' \ dx \]

OpenStudy (anonymous):

yes

OpenStudy (jamesj):

We want to 'get rid' of the ln x because it's tricky. Hence it's much better here to start with u' = x and v=ln x

OpenStudy (anonymous):

ok

myininaya (myininaya):

do you know how to do \[\int\limits_{}^{}ue^u du\]

OpenStudy (anonymous):

probably not

myininaya (myininaya):

Oh i was going to say to could have put it in this form with a substitution nvm

OpenStudy (anonymous):

this problem in in the intragtion by parts section

OpenStudy (jamesj):

ok. with the suggested u' and v, what have you got?

myininaya (myininaya):

You can use substitution +integration parts if you wanted

OpenStudy (jamesj):

@myin: it really is easy to do by parts. I don't think it's helpful to substitute here.

myininaya (myininaya):

I agree with you James. I was just trying to see if that form was easier for him to integrate by parts.

OpenStudy (anonymous):

1/x - integral of

OpenStudy (anonymous):

is that the right start?

OpenStudy (jamesj):

If u' = x, what is u? And with v=ln x, what is v'? u = .... v' = ...

OpenStudy (anonymous):

u= 1/2x^2

OpenStudy (anonymous):

v = 1/x

OpenStudy (jamesj):

u = x^2/2, yes v = 1/x, yes. hence the integral equals \[ \frac{x^2 \ln x}{2} - \int \frac{x^2}{2x} \ dx \]

OpenStudy (jamesj):

Now, the new integral simplifies and you can integrate it directly.

OpenStudy (jamesj):

v' = 1/x

OpenStudy (anonymous):

thanks for your help, im just terrible at math, i just cant grasp this stuff

OpenStudy (anonymous):

i get lost when ur splitting the v and v 1's

OpenStudy (anonymous):

which is which

OpenStudy (jamesj):

u' = x, u = x^2/2 v = ln x, v' = 1/x Hence \[ \int u'v = uv - \int uv' \] \[ \int x \ln x \ dx \ = \ \frac{1}{2}x^2 \ln x - \int \frac{x^2}{2} \frac{1}{x} \ dx \] yes?

OpenStudy (anonymous):

ok

OpenStudy (jamesj):

So what's your final answer?

OpenStudy (anonymous):

james when ur done here can you jump in my post plz?